HIGH RESOLUTION ELECTRONIC SPECTROSCOPY OF 9,10-DIHYDROPHENANTHRENE ^{a}. STUDY OF THE INVERSION MOTION OF A FLOPPY MOLECULE IN THE GAS PHASE ^{b}

\underline{LEONARDO ALVAREZ}, and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

Rotationally resolved S_{1} ← S_{0} fluorescence excitation spectra of 9,10-dihydrophenanthrene (DHPH) have been studied in a collision-free molecular beam environment. DHPH exhibits two well defined Franck-Condon progressions of vibronic bands. From the analyses of its fully resolved electronic spectra, the first progression has been assigned to the out of plane CH_{2}-CH_{2} bridge deformation inversion motion of the molecule. The second progression, split into doublets of alternating intensities, is likely to be caused by combination of vibrational modes upon excitation. Detailed information about the assignments and inversion barriers in both electronic states will be discussed.

aM. Z. Zgierski, F. Zerbetto, Y. Shin, and E. C. Lim, J. Chem. Phys. \underline{96 (10)}, 1992, pp 7229.

bWork supported by NSF